what volume (in ml) of water would have to be added to 150.0ml of 0.450 m potassium chloride, kcl, to give a solution with a concentration of 0.100 m?
1
Expert's answer
2020-03-30T10:51:32-0400

Solution.

"n(KCl) = 0.15 \times 0.450 = 0.0675 \ mol"

"V'(KCl) = \frac{n(KCl)}{C'(KCl)}"

V'(KCl) = 0.675 L = 675 ml

"V(H2O) = 675 - 150 = 525 ml"

Answer:

V(H2O) = 525 ml